OpenStudy (anonymous):

y varies inversely as three times x. When x = 6, y = 4. Find x when y = 8

OpenStudy (paxpolaris):

\[\Large y \propto \frac 1 {3x}\] \[\Large \implies 3xy=k\] so, k = 3*6*4 3*6*4= 3*x*8

OpenStudy (paxpolaris):

as a side note: saying “y varies inversely as 3x” also means “y varies inversely as x”